OpenStudy (anonymous):

Is x = 1 a solution of the equation 2 – 8x = –6?

OpenStudy (anonymous):

Plug it in, and find out, plug it in for x.

OpenStudy (anonymous):

I would if I had a claculator. sadly, I don't.

OpenStudy (anonymous):

Ok, let us work on this manually. :D

OpenStudy (anonymous):

8(1)

OpenStudy (anonymous):

What does that equal?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Ok, so the equation we have is 2-8=? if you multiply with a lower number before a bgger number, your answer will obviously be negative.

OpenStudy (anonymous):

okay, so is that what I am to do?

OpenStudy (anonymous):

Yep! Have u got it?

OpenStudy (anonymous):

no, because my initial equation is 2 - 8x = -6 I don't understand where you got your equation from...

OpenStudy (anonymous):

That's what I did... 2 – 8x = –6 2 - 8(1)= -6 And you solve for 8(1)=8 Then u solve the equation.

OpenStudy (anonymous):

2-8=?

OpenStudy (anonymous):

ohhhhh okay! i get it! THANK YOU!

OpenStudy (anonymous):

You r welcome! Can u solve it on ur own now?

OpenStudy (anonymous):

yes!

OpenStudy (skullpatrol):

2 – 8x = –6 2 - 8(1)=?= -6 -6= -6

OpenStudy (anonymous):

Okay, I will lave you to it! :D good luck. :) 

OpenStudy (anonymous):

skullpatrol your answer is wrong.

OpenStudy (skullpatrol):

Why?

OpenStudy (anonymous):

Nope, he is right... 2-8= -6, so that makes your question true.